OpenStudy (anonymous):

what is the slope of a line that crosses through the points (-4,-5) and (6,8)

OpenStudy (anonymous):

Slope is:\[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

x1: -4 x2: 6 y1: -5 y2: 8 Now fill in the values to solve.

OpenStudy (anonymous):

i dont understand it isnt the formula y+mx+b

OpenStudy (anonymous):

I think that there may be more than one formula, but this one works as well. \[\frac{ 8-(-5) }{ 6-(-4) }\]This is just the same thing as I said before, but I put the values in for you.

OpenStudy (anonymous):

after that do we square it and and then square root???

OpenStudy (anonymous):

No, just solve the equation on the top and bottom of the fraction bar. I'll do it for you this time, but I want you to let me know if you don't understand. 8-(-5)=13 6-(-4)=10 Then we divide them. 13/10=1.3, or: \[1\frac{ 1 }{ 3 }\]So, the slope if the line is 1.3 or 1 1/3. Let me know if you need further help.

OpenStudy (anonymous):

ohh i understand now thanks(:

OpenStudy (anonymous):

your welcome.